# RPNT: Robust Pre-trained Neural Transformer - A Pathway for Generalized Motor Decoding

### v1

@@ 4.4 Ablation Studies
We conducted ablation studies on T-RT task and the novel B-CS task in the FS-SFT regime to evaluate the contribution of novel components of position encoding, attention mechanism, and masking strategy to the performance of RPNT and report the results in Tables 44 (see appendix table 8 and 8 for the mean $\pm$ std), and 4. Additional sweeping analyses, e.g., transformer layer, attention head, and context-based kernel size, are included in the appendix section K. MRoPE. We compared MRoPE against three alternative positional encoding approaches: sinusoidal PE, standard RoPE, and multi-dimensional learnable PE. In Table 4, we compare MRoPE against alternative positional encodings and observe that MRoPE achieves the highest $R^{2}$ score. Improvement of $\approx 3\%$ over standard RoPE validates the extension of RoPE to multiple dimensions Context-based Attention Mechanism. Table 4 demonstrated the importance of context-dependent attention design. Replacing context-based attention with standard self-attention resulted in a substantial performance drop of $\approx 5\%$. This significant gap highlights the effectiveness of the adaptive kernel generation mechanism in handling the non-stationary nature of neural recordings (see appendix table 6 for kernel size studies). Random Uniform Masking Strategy. Fixed masking ratios, whether symmetric or asymmetric across neuron and temporal dimensions, consistently underperformed the RPNT random uniform masking approach (Table 4).
5 Discussion and Conclusion
We present RPNT, a robust pretrained neural transformer comprised of three novel components (MRoPE, context-based attention mechanism, and uniform random masking strategy). Pretraining RPNT with SSL objectives eliminates the need for behavioral labels and allows for finetuning it in a few-shot regime (FS-SFT) in decoding downstream tasks (Gunel et al., 2020; Ding et al., 2023). Our experiments show that RPNT outperforms in generalization to a variety of downstream decoding tasks. These advances pose RPNT as a promising pathway toward neural foundation modeling for neural decoding and brain-computer interfaces. Our work also has limitations. First, our modeling was limited to cortical motor areas, excluding brain regions such as the frontal cortex and the basal ganglia, which play a key role in motor control (Trautmann et al., 2025). Future work should incorporate recordings from these regions to pretrain more comprehensive neural population models. Second, center-out and random reaching tasks represent simplified and stereotyped motor behaviors. Natural behaviors and cognition tend to be more complex across species and even include feedback interactions with the environment. Validating RPNT on such datasets will be the next step (Pei et al., 2021; Willett et al., 2023; Karpowicz et al., 2024; de Vries et al., 2023; Safaie et al., 2023; Banga et al., 2025). Last, RPNT was developed for a single modality. Developing multimodal RPNT will be the next work (Wang et al., 2024a; Zhang et al., 2025).
Acknowledgments and Disclosure of Funding
@@ 5 Discussion and Conclusion
The authors acknowledge the partial support of the Harnessing the Data Revolution (HDR) Institute: Accelerated AI Algorithms for Data-Driven Discovery (A3D3), National Science Foundation grant (PHY-2117997 A.L.O., E. S., H.F.). The authors thank the support of the Departments of Electrical and Computer Engineering, Applied Mathematics, and Bioengineering. For the neuropixel dataset collection, it was supported in part by a National Center for Advancing Translational Sciences of the National Institutes of Health fellowship (TL1 TR002318, R.A.C.), a Nakajima Foundation fellowship (T.O.), a Simons Collaboration for the Global Brain Pilot award (898220, A.L.O.), the National Institute of Neurological Disorders and Stroke (NIH R01 NS134634 and NIH UF1NS126485, A.L.O.), and the NIH Office of Research Infrastructure Programs (P51 OD010425).
Appendices
Appendix A Ethics statement
@@
Our work focuses on algorithm design and evaluation for neural data. To evaluate the RPNT model, we used two datasets: a public dataset and a neuropixels dataset. The public dataset is from Perich et al. (2018). Data privacy and ethics protocols for experimental procedures related to neural data collection for this data were addressed in the original publication. The neuropixels dataset was collected at the University Institution and the National Primate Research Center Canfield et al. (2025). All procedures were conducted in compliance with the NIH Guide for the Care and Use of Laboratory Animals and were approved by the institutional Animal Care and Use Committee.
Appendix B The use of Large language models (LLMs)
We used the large language model GPT for checking grammar and spelling, and improving the clarity of our manuscript’s language. All scientific content, experimental design, analysis, and interpretations are entirely the original work of the authors.
Appendix C Reproducibility statement
@@
All hyperparameters, architectural details, and training configurations necessary for reproducing our results are provided in Appendix O. Data preprocessing steps and experimental protocols are described in Appendix N. Code will be made publicly available on GitHub upon paper acceptance.
Appendix D Broader Impact
Our results have a broad impact on future brain-computer interface (BCI) designs. Current BCI decoders typically process neural activity signals in isolation, missing the rich information encoded in historical recordings (Le et al., 2025). Our RPNT’s superior performance on held-out cases demonstrates that the pretraining strategy can substantially improve decoding robustness (Chen and others, 2025). Second, the computational visualization framework provides a data-driven tool to investigate the spatial task information distribution (Canfield et al., 2025; Fang and others, 2024a) and potential directional information flow across brain sites. Third, while demonstrated on the motor cortex, our RPNT framework applies to broad neural decodings for wide BCI applications such as closed-loop neuromodulation systems (Fang and Yang, 2021; 2023; Fang and others, 2024b) and lightweight neural decoders (Fang and others, 2025).
Appendix E Data Pre-processing Pipeline

### v2

@@ Results
Discussion and Conclusions
We present RPNT, a robust pretrained neural transformer designed for generalized motor decoding across sessions, subjects, behavior types, and recording sites. RPNT combines three novel components — multidimensional rotary positional embedding (MRoPE), context-based attention mechanism, and uniform random masking strategy – with a self-supervised learning (SSL) framework that eliminates the need for behavioral labels during pretraining. Unlike existing pretrained neural decoders such as POYO and its subsequent work [1, 2, 43], which rely on supervised objectives that require behavior labels during pretraining, RPNT learns robust neural representations purely from spike activity and adapts to downstream tasks via few shot supervised finetuning (FS-SFT). We validated RPNT on two datasets: a multi-session, multi-subject, multi-task microelectrode benchmark [39] (LTRCH) and a new multi-site Neuropixel 1.0 dataset [4] comprising 17 recording sites spanning large portions of PMd and M1 over approximately one year (NPCS). The NPCS dataset evaluated in this work enabled the first cross site generalization benchmark for pretrained neural decoders. Across all four evaluation scenarios, RPNT consistently surpassed existing state-of-the-art models, notably outperforming supervised pretrained baselines that had access to behavior labels throughout pretraining. These results, together with comprehensive ablation studies, establish RPNT as a promising pathway toward neural foundation models for brain decoding.
The algorithmic innovations in RPNT directly address key limitations in existing neural transformer architectures. First, standard positional encodings, such as sinusoidal [55] and learnable [15], cannot represent the multi-dimensional experimental metadata inherent to neural recordings, such as electrode coordinates, session identities, and behavior types. MRoPE extends rotary positional embedding [51] to arbitrary dimensions by partitioning the model dimension into groups that each encode distinct metadata axes. This design draws on recent extensions of multi-dimensional RoPE in vision language models [58], adapting them to the structure of neural recording configurations. Our ablation confirmed an approximate $3\%$ improvement over standard RoPE, validating the benefit of metadata-aware positional encoding for cross configuration generalization. Given the many sources of potential variability in neural data, even beyond those explored in our work (e.g., neural state variations like alertness/attention, measurement modalities), our results suggest that metadata-aware positional encoding strategies will be important components for neural foundation models and robust decoding architectures.
Second, neural signals exhibit pronounced non-stationarity due to drift across sessions [9], yet standard self-attention primarily captures global neural dynamics. Our context-based attention mechanism introduces learnable convolution kernels [25] applied to global attention scores to capture local temporal structure, drawing on principles from non-stationary time-series modeling [32]. Our ablation analyses revealed a substantial performance gap when replacing context-based attention with standard attention, approximately $5\%$ on cross-task and up to $16\%$ on the cross-site evaluation. This result indicates that the proposed attention mechanism is critical for decoding under distributional shifts present in neural data.
Third, prior neural transformers employ bidirectional masked autoencoding [64, 63, 65, 26] for offline denoising, which violates the causal structure required for real-time applications. Our causal masked neural modeling enforces unidirectional information flow, transforming the objective from interpolation to extrapolation. Combined with uniform random masking sampled from $\mathcal{U}(0,1)$, this strategy removes the need for a masking ratio hyperparameter and exposes the model to a range of reconstruction difficulties, yielding more robust representations than any fixed masking configuration. Collectively, these three components form an integrated design that addresses spatial heterogeneity, temporal non-stationarity, and causal representation learning in neural data.
Our work has broad implications for brain-computer interface (BCI) design. Current BCI decoders typically process neural activity in isolation and require frequent recalibration to maintain performance across sessions [9, 30, 11, 49, 21, 14, 35]. RPNT demonstrated superior cross session and cross site generalization, which shows that self-supervised pretraining on unlabeled neural data can substantially improve decoding robustness [31, 67, 63]. More robust decoding architectures may help reduce the calibration burden that remains a major obstacle to long-term BCI deployment [42, 29]. Practically, the RPNT pretraining-finetuning paradigm enables a scalable BCI workflow, where a model is pretrained once on large-scale neural recordings and then rapidly adapted to individual users or sessions via FS-SFT with minimal labeled data [7]. Such data efficiency is especially relevant in clinical settings where labeled trial collection is time-consuming and burdensome for patients. Moreover, while demonstrated on motor cortex velocity decoding, the RPNT architecture is not restricted to a specific decoded variable or brain region. The framework has natural extensions to broader applications, including closed-loop neuromodulation systems [19, 20, 17] and lightweight neural decoders for resource-constrained clinical settings [18].
While promising, the limitations of RPNT and our evaluation of it merit further future work. Our modeling and evaluation were restricted to cortical motor areas (PMd and M1), excluding additional cortical and sub-cortical brain regions, such as the frontal cortex and basal ganglia, which contribute to motor control and may have distinct dynamics [53]. Future work should incorporate recordings from these regions to pretrain more comprehensive neural population models that capture the distributed nature of motor processing. Second, center-out and random target reaching tasks represent relatively simplified and stereotyped motor behaviors. Natural behaviors, such as continuous object manipulation, sequential movements, and cognitive tasks, involve richer dynamics and feedback interactions with the environment. Validating RPNT on more naturalistic behavioral datasets will be essential to establish broader applicability [38, 59, 29, 13, 44, 3]. Third, each RPNT model instance currently processes a single neural spike modality. Brain function involves coordinated dynamics across multiple signal types; local field potentials (LFP), for instance, carry mesoscale oscillatory information complementary to single-unit spiking. Extending RPNT to a multimodal framework that jointly models heterogeneous neural signals [57, 68, 5] represents a key next step toward more comprehensive neural foundation models.
Methods
We describe the technical design of RPNT in this section. Briefly speaking, RPNT contains the following three innovative considerations that emerged from the prior investigation of neural transformer architectures [64, 63, 30, 65, 43, 1]. Figure 4 illustrates the overall architecture design. It includes: 1) Multidimensional rotary positional embedding (MRoPE) that aggregates experimental metadata such as site coordinates, session ids, and behavior types; 2) A context-based attention mechanism that applies convolution kernels to global attention to learn local temporal structure and handle the non-stationarity of neural activity; 3) A robust self supervised learning (SSL) objective with stochastic causal masking strategies and contrastive representation learning. For following each section, we first describe the major difference from the prior work and then the design details.
MRoPE: Multi-dimensional Rotary Positional Embedding for neural session configurations
While standard sinusoidal encodings [55] and learned positional embeddings [15] are generally applicable, positional encoding remains challenging for neural data because of signal drift across sessions [9]. Recently, rotary positional embedding (RoPE) [51, 1] has become a strong positional encoding approach in pretrained transformer models. Extensions of RoPE to multimodal settings have also shown promise in handling complex positional relationships beyond simple sequential ordering. For example, Qwen2 [58] demonstrates an effective multi-dimensional RoPE design that captures both sequence position and modality specific structure. Motivated by these developments, we argue that positional encoding for neural data should account for joint recording configurations beyond time alone, so that each neural recording session configuration is represented explicitly through the transformer positional embedding. For example, in recordings across multiple brain sites, an effective positional embedding should model both spatial site location and temporal relationships in order to generalize to unseen site configurations. Thus, we propose a novel extension to RoPE, extending it to multiple dimensions, MRoPE, where $M$ denotes the number of configuration dimensions used to aggregate experimental metadata. This design is more appropriate for neural recordings and fundamentally different from Qwen2, developed for multimodal interactions.
Using the "Neuro Pixels across Cortical Sites“ dataset (NPCS) as an example, we define MRoPE below. The minimal dimension for this dataset $M$ is $M=3D$, where $D$ stands for the number of electrodes and $3$ corresponds to two dimensional spatial coordinates (denoted by $(x_{s},y_{s})$) and time $t$. To construct the rotational matrix, we partition the model dimension ($d_{\text{model}}$) into three groups independently ($d_{\text{coord}}=\frac{d_{\text{model}}}{3}$) to represent dimensions each for $x$-coordinate, $y$-coordinate, and temporal position. For each group, we define rotation frequencies:
where $i\in[0,d_{\text{coord}}/2)$. Spatial dimensions use a lower frequency ($f=5000$) to preserve coherence across nearby recording sites, whereas the temporal dimension uses the standard RoPE frequency ($f=10000$) for fine-grained temporal resolution.
Rotation operation. In contrast to common positional encodings that are added to embeddings, MRoPE, similarly to RoPE in mathematical formulation, operates by applying rotations directly to query and key vectors during attention computation. For a 3D vector that correponds to $M=3$, $(x_{s},y_{s},t)$, we construct a block-diagonal rotation matrix $\mathbf{R}_{\text{3D}}(x_{s},y_{s},t)$(see (2)) that independently rotates each pair of dimensions. The attention score between locations $i$ and $j$ follows $\mathbf{q}_{i}^{T}\mathbf{k}_{j}=\mathbf{q}^{T}\mathbf{R}_{\text{3D}}^{T}(x_{i},y_{i},t_{i})\mathbf{R}_{\text{3D}}(x_{j},y_{j},t_{j})\mathbf{k}=f(x_{j}-x_{i},y_{j}-y_{i},t_{j}-t_{i})$, which preserves the key property of relative position encoding. This operation supports zero shot generalization to arbitrary brain site configurations while preserving rotational invariance (also see Appendix 3).
MRoPE is designed for arbitrary dimensions, allowing it to represent recording configurations that vary across subjects, recording times, locations, and behavior types (see Figure 4). Specifically, we partition the model dimension $d_{\text{model}}$ into $M$ groups according to the number of the recording configuration dimensions, i.e., $\theta^{(1)}_{i}=\frac{1}{f_{1}^{2i/d_{1}}},\quad\cdots\quad,\theta^{(\mathcal{M})}_{i}=\frac{1}{f_{M}^{2i/d_{M}}}$, where $d_{1},\cdots,d_{M}$ are the dimensions of each group and together sum to $d_{\text{model}}$. The $f_{1},\cdots,f_{M}$ are the corresponding frequencies. Indeed, for the LTRCH benchmark, we use MRoPE with $M=4D$ to represent behavior type, subject identity, recording time, and temporal position. We used an equal dimension partition ($d_{m}=\frac{d_{model}}{4}$) and set frequencies with 10, 100, 1000, and 10000, respectively (see Appendix 3 for further details).
Figure 4: A schematic of the components of RPNT. Components shown in black indicate the standard transformer signal flow, including no masking and a standard attention mechanism. The proposed components are MRoPE (green), context based attention (cyan), and the uniform random masking strategy (pink).MRoPE incorporates experimental metadata into the transformer positional embeddings. The context-based attention captures local temporal structures. The input spike activity is randomly masked according to a uniform distribution for causal masked neural modeling. The masked spike activity is then projected into neural embeddings, forming query, key, and value representations. MRoPE is applied to produce rotary queries and keys. Attention scores are modulated by context based convolutional kernels and combined with causal masking before aggregation with the values.
Context-based attention mechanism
Adaptive attention mechanisms have been proposed for domains with nonstationary data, such as video [8]. These mechanisms are based on learnable kernels to modulate the attention based on local context. Further, a sliding window attention was introduced to capture local structure [72], while a hierarchical attention across multiple timescales was employed [62]. Non-stationary Transformers [32] introduce series stationary and nonstationary attention to handle distribution shifts in time series forecasting. FEDformer [71] employs frequency-enhanced decomposed architectures to capture both seasonal and trend patterns in nonstationary signals. Crossformer [70] utilizes dimension-segment-wise attention to capture cross-time and cross-dimension dependencies. Motivated by these works and nonstationary nature of neural activity [9], we introduce a context-based attention mechanism via learnable convolution kernels [25] applied to global attention scores to capture local temporal structure. This mechanism incorporates historical context and attention pooling into the attention computation. We also illustrate the attention steps (cyan color) in Figure 4.
Context Generation. Given a batch neural input $\mathbf{x}\in\mathbb{R}^{B\times T\times D}$ and historical data $\mathbf{H}_{\text{hist}}\in\mathbb{R}^{B\times T_{\text{hist}}\times D}$ containing only past timesteps, we generate a context vector through attention pooling:
We note that $T_{\text{hist}}\subseteq T$. In practice, a simple choice for $T_{\text{hist}}$ is $T_{\text{hist}}=T$, which uses the available historical data to generate the convolution kernels. A more computationally efficient alternative is to use a receding window, $T_{\text{recede}}$, e.g., the most recent $K$ time bins. In this work, we use the first option due to its simplicity.
Dynamic kernel generation. We parameterize the context vector using MLP layers to generate 2D convolution kernels for each attention head
where $h\in[1,H]$ indexes the attention heads and $(K_{1},K_{2})$ are hyperparameters defining the kernel dimensions.
Kernel-enhanced attention. Finally, we apply a 2D convolution kernel after calculating the standard attention matrix to further capture the local attention structure
where $\mathbf{M}_{\text{causal}}$ is the causal masking ($\mathbf{M}_{\text{causal}}[i,j]=\mathbb{1}[j\leq i]$). This design allows each attention head to learn specialized temporal dependencies for handling neural nonstationarity (see Figure 4). We follow the convention to apply the causal masking at the very end of the attention operation to preserve the autoregressive property. An alternative causal masking implementation can also be applied within the convolutional operation. In Appendix 7, we show that both causal masking implementations make no difference. Thus, we followed the first implementation due to it simplicity.
Neural transformer architecture
Neural encoder. With the above two components, we construct RPNT’s neural transformer encoder, which decouples temporal (for both the NPCS and the LTRCH datasets) and cross site processing (only for the NPCS dataset). First, we build a temporal encoder. Given neural spike activity $\mathbf{X}\in\mathbb{R}^{B\times S\times T\times N}$ (batch, sites, time, neurons) and corresponding experimental metadata, we first embed each recording site $\mathbf{H}_{s}^{(0)}=\text{Linear}(\mathbf{X}_{:,s,:,:})$. Then, we build a temporal encoder that consists of $L$ transformer layers with MRoPE and context-based attention
Most neural representation learning and neural population modeling can be handled by this temporal encoder alone by setting the number of recording sites $S$ to $S=1$, corresponding to repeated neural measurements across sessions. When modeling cross site neural dynamics is needed, such as in sampled Neuropixel recordings for cross site activity (NPCS dataset), we further concatenate a spatial encoder after the temporal encoder. The spatial encoder models cross site interactions through time stepwise processing while naturally preserving causality. At each timestep $t$, we follow the convention to apply the standard multi-head attention (MHA)
$\mathbf{Z}_{:,:,t,:}^{(l+1)}=\mathbf{Z}_{:,:,t}^{(l)}+\text{MHA}(\text{LN}(\mathbf{Z}_{:,:,t,:}^{(l)})),$ (7)
where $\mathbf{Z}^{(0)}=\mathbf{H}^{(L)}$ is taken from the final layer of the temporal encoder. We especially note that the resulting MHA weights can provide interpretable cross-site functional connectivity for NPCS dataset. In our experiments on LTRCH microelectrode benchmark [39], we focus on demonstrating and validating the temporal encoder. Further, we use the NPCS Neuropixel dataset to provide extra data-driven insights into the MHA map, which discovers the potential brain site interaction.
Neural decoder. Our main goal in RPNT is to obtain a pretrained transformer encoder that can later be finetuned for downstream tasks or frozen for data-driven functional connectivity visualization. Accordingly, we design the decoder to be lightweight such that $\text{Decoder}_{\theta}:\mathbb{R}^{B\times T\times D}\rightarrow\mathbb{R}^{B\times T\times N}$ maps encoded representations to Poisson rate parameters through two feedforward blocks and an output projection:
$\boldsymbol{\lambda}$ $=\text{Softplus}(\text{Linear}(\mathbf{Z}))$ (9)
Causal masked neural modeling
With the neural transformer setup, we next introduce a pretraining framework tailored for neural spike data that robustly performs self-supervised learning (SSL) using masked causal Poisson reconstruction loss with a stochastic masking strategy (uniform random distribution) and contrastive loss for cross-site latent representations.
Uniform random masking strategy. Unlike fixed masking strategies, which require careful tuning in prior work [50, 65, 69, 26], we apply masking in both space (neurons) and time according to a stochastic uniform distribution such that $p_{m}\sim\mathcal{U}(0,1)$, at each batch. This choice is motivated by prior supervised motor decoding studies [1, 30]. In POYO [1], dynamic dropout was shown to improve decoding performance, and in SPINT [30], a uniform neuron sampling strategy further improved supervised motor decoding. These findings motivated us to extend stochastic masking to both temporal and neuronal dimensions during SSL pretraining. This stochastic strategy removes the masking rate hyperparameter while ensuring that the model encounters diverse reconstruction difficulties during pretraining, i.e., $\mathbf{M}_{:,t,n}\sim\text{Bernoulli}(1-p_{m})$.
Self-supervised learning objective. Distinct from existing denoising-based transformer work [64], our approach maintains causality, i.e., reconstruction at time $t$ depends only on unmasked inputs from $t^{\prime}\leq t$. This adapts the standard masked autoencoding objective [26] into a predictive task aligned with the autoregressive nature of neural mechanisms (see discussion section). For masked position $(t,n)$, RPNT reconstructs it with respect to
where $\hat{\lambda}_{:,t,n}=\text{Decoder}(\mathbf{Z}_{:,t,d})$ and $\epsilon=10^{-8}$ provides numerical stability. We further include an auxiliary objective to encourage site invariant representations (see Appendix 4 for details). To this end, the total combined pretraining loss function $\mathcal{L}=\mathcal{L}_{\text{recon}}+\mu\mathcal{L}_{\text{contrast}}$ balances reconstruction objective with robust site-invariant representation.
Downstream evaluation framework
$\hat{\mathbf{y}}_{t}=\text{MLP}_{\text{task}}(\mathbf{H}_{:,t}),$ (12)
where $\mathbf{H}_{:,t}$ is the temporal encoder outputs across at time $t$. We use mean-squared error (MSE) for SFT on velocity regression tasks and evaluate using $R^{2}$ metric. We additionally visualize functional connectivity based on the spatial attention maps across time. The spatial encoder’s cross site attention weights directly encode functional relationships. We extract the connectivity matrix $\mathbf{C}\in\mathbb{R}^{T\times S\times S}$ by averaging attention weights across layers, i.e., $\mathbf{C}_{i\rightarrow j}(\text{attention})=\frac{1}{L}\sum_{l=1}^{L}\text{Attention}_{i\rightarrow j}^{(l)}$. The resulting visualizations are provided in Appendix 8.
Acknowledgements
@@
The authors acknowledge the partial support of the Harnessing the Data Revolution (HDR) Institute: Accelerated AI Algorithms for Data-Driven Discovery (A3D3), National Science Foundation grant (PHY-2117997 A.L.O., E. S., H.F.). The authors thank the support of the Departments of Electrical and Computer Engineering, Applied Mathematics, and Bioengineering at the University of Washington. For the neuropixel dataset collection, it was supported in part by a National Center for Advancing Translational Sciences of the National Institutes of Health fellowship (TL1 TR002318, R.A.C.), a Nakajima Foundation fellowship (T.O.), a Simons Collaboration for the Global Brain Pilot award (898220, A.L.O.), the National Institute of Neurological Disorders and Stroke (NIH R01 NS134634 and NIH UF1NS126485, A.L.O.), and the NIH Office of Research Infrastructure Programs (P51 OD010425).
Author contributions statement
A.L.O. and E.S. supervised the project. H.F., A.L.O., and E.S. conceived the RNPT framework. H.F. implemented the code. HF wrote the initial version of the manuscript. H.F., A.L.O., and E.S. discussed and analyzed the results. H.F., A.L.O., and E.S. reviewed and revised the manuscript. R.A.C, T.O., and B.M. conducted the Neuropixel data collection experiments. All authors reviewed the manuscript before submission.
Ethics statement
@@
Our work focuses on algorithm design and evaluation for neural data. To evaluate the RPNT model, we used two datasets: a public dataset and a neuropixels dataset. The public dataset was introduced in the prior work [39] and then made available online [40]. Data privacy and ethics protocols for experimental procedures related to neural data collection for this data were addressed in the original publication. The neuropixels dataset was collected at the University of Washington and the Washington National Primate Research Center [4]. All procedures were conducted in compliance with the NIH Guide for the Care and Use of Laboratory Animals and were approved by the institutional Animal Care and Use Committee.
Reproducibility statement
@@
All hyperparameters, architectural details, and training configurations necessary for reproducing our results are provided in Appendix 9. Data preprocessing steps and experimental protocols are described in Appendix 2. The code will be made publicly available on GitHub upon acceptance of the paper.
Appendix
1 Micro-electrode Data Pre-processing Pipeline